in three moles of ethane calculateNumber of molecule of ethane

Grade:11

1 Answers

Asmit Jain
6 years ago
18.069*1023 molecules
because no of molecules=no. of moles * avogadro no.
since n=3 NA=6.023*1023
Therefore, no. of moleculed=n*NA=3*6.023*1023=18.069*1023 molecules.